# Towards Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v2

@@ 5 Logit normalisation improves the SC of LS-trained models
$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=\boldsymbol{v}/\left\lVert\boldsymbol{v}\right\rVert_{p}=\boldsymbol{v}\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
@@
Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. A.2.
Figure 8: Visualisation of the effect of logit normalisation for segmentation (same scene as Fig. 1). Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model. We provide more figures in Appendix F.
@@
Although Cattelan & Silva (2024) empirically validate this approach on a large number of pre-trained models, it remains unclear as to why it is so effective, or why it isn’t effective sometimes (further discussion in Sec. E.5). Let us shed light on the interaction between logit normalisation and LS. If we examine Eq. 16, we see that it resembles the softmax, however, the logits are raised to power $p$ rather than exponentiated. Crucially, whilst the softmax is invariant to uniform shifts in the logits,
@@
this is not the case for $\boldsymbol{v}^{\prime}$. In fact, we have the following inequality for positive [^6] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits, meaning the behaviour discussed still occurs empirically (Secs. E.3 and 20). logits $\boldsymbol{v}$:
@@
The proof can be found in Appendix D. Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of the normalised max logit $v^{\prime}_{\text{max}}$. That is to say logit normalisation increases uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$. Recall how in Sec. 4.2 we show how imbalanced logit regularisation leads to degradation in SC. In particular Fig. 5 shows how LS leads to higher $v_{\text{max}}$ given $\pi_{\text{max}}$ for errors ✗ – this implies that, independent of the value of $\pi_{\text{max}}$, information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$. We see now that logit normalisation will increase the uncertainty (ranking) of errors ✗  relative to correct predictions ✓  using the information in $v_{\text{max}}$, leading to improved SC for an LS model. That is to say logit normalisation effectively reverses the effect of the imbalanced logit regularisation from LS, improving SC. We note that Cattelan & Silva (2024) find that models that are less confident (on both ✓ and ✗) tend to benefit more from logit normalisation, again pointing towards LS. Given our analysis and empirical results, we strongly recommend logit normalisation for LS-trained models when performing SC. On the other hand, Fig. 5 also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE model. This explains why for CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.
@@ 6 Related work
Prediction with rejection. Selective classification falls into the broader problem setting of prediction with rejection. In the case of SC, misclassifications are to be rejected (El-Yaniv & Wiener, 2010). The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023). Another scenario is out-of-distribution (OOD) detection (Yang et al., 2021; Hendrycks & Gimpel, 2017), where data from outside of the training distribution are to be rejected. There is a plethora of research in this field (Xia & Bouganis, 2022a; Sun et al., 2021; Liu et al., 2020; Zhang et al., 2023; Liu et al., 2023; Wang et al., 2022; Hendrycks et al., 2022). Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data (Jaeger et al., 2023; Xia & Bouganis, 2022b; Kim et al., 2023). Notably, Deep Ensembles (Lakshminarayanan et al., 2020) have arisen as a reliable method to improve performance in all three scenarios (Kim et al., 2023; Xia & Bouganis, 2023; Laurent et al., 2023; 2024). We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
@@ 7 Concluding remarks
In this work, we elucidate the effect of label smoothing (LS) on selective classification (SC). Our experiments across various tasks and architectures show that LS leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves accuracy. By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS regularises/suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions. We then investigate post-hoc logit normalisation as a method to improve the degraded SC performance caused by LS. We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher. We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
@@
To ensure transparency, we use publicly available datasets: CIFAR-100, ImageNet, and Cityscapes. Our detailed experimental methods are outlined in Appendix B. Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix D.
To help replicate our work, we will soon share our source code on GitHub, including configuration files and the training code. We also plan to release our most important models on Hugging Face.
@@
Appendix A Additional results
A.1 Illustrative examples of ImageNet overconfidence
@@
A.2 Complete results for ViT and DeepLabV3+
@@
  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix F.
Figure 10: RC curves when evaluating on the combination of ImageNet and ImageNet-sketch. We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced. This shows that LS leads to increasing overconfidence on less-well-fit data.
@@
CE #samples 8994 1094
#errors 93 80
error rate 1.0 7.3
LS $\alpha=0.1$samples 8561 1527
#errors 187 211
error rate 2.2 13.8
LS $\alpha=0.2$samples 8171 1917
#errors 244 365
error rate 3.0 19.0
LS $\alpha=0.3$samples 7915 2173
#errors 297 541
error rate 3.8 24.9
@@
CE #samples 9528 560
#errors 88 43
error rate 0.9 7.7
LS $\alpha=0.1$samples 9332 756
#errors 130 83
error rate 1.4 11.0
LS $\alpha=0.2$samples 9117 971
#errors 177 164
error rate 1.9 16.9
LS $\alpha=0.3$samples 8966 1122
#errors 208 195
error rate 2.3 17.4
@@
A.3 Small-scale CIFAR experiments
@@
Appendix B Reproducibility
We recall that all the code used in this project will be made available soon to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix A. All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
B.1 Image classification

### v3

@@ 5 Logit normalisation improves the SC of LS-trained models
$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=[\boldsymbol{v}+s]/\left\lVert\boldsymbol{v}+s\right\rVert_{p}=[\boldsymbol{v}+s]\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}+s|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
@@
Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. B.2.
Figure 8: Visualisation of the effect of logit normalisation for segmentation (same scene as Fig. 1). Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model. We provide more figures in Appendix G.
@@
Although Cattelan & Silva (2024) empirically validate this approach on a large number of pre-trained models, it remains unclear as to why it is so effective, or why it isn’t effective sometimes (further discussion in Sec. F.4). Let us shed light on the interaction between logit normalisation and LS. If we examine Eq. 16, we see that it resembles the softmax, however, the logits are raised to power $p$ rather than exponentiated. Crucially, whilst the softmax is invariant to uniform shifts in the logits,
@@
this is not the case for $\boldsymbol{v}^{\prime}$. In fact, we have the following inequality for positive [^6] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits, meaning the behaviour discussed still occurs empirically (Secs. F.3 and 22). logits $\boldsymbol{v}$:
@@
The proof can be found in Appendix E. Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of the normalised max logit $v^{\prime}_{\text{max}}$. That is to say logit normalisation increases uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$.
Recall how in Sec. 4.2 we show how imbalanced logit suppression leads to degradation in SC. In particular Fig. 5 shows how LS leads to higher $v_{\text{max}}$ given $\pi_{\text{max}}$ for errors ✗ – this implies that, independent of the value of $\pi_{\text{max}}$, information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$. We see now that logit normalisation will increase the uncertainty (ranking) of errors ✗  relative to correct predictions ✓  using the information in $v_{\text{max}}$, leading to improved SC for an LS model. That is to say logit normalisation effectively reverses the effect of the imbalanced logit suppression from LS, improving SC. We note that Cattelan & Silva (2024) find that models that are less confident (on both ✓ and ✗) tend to benefit more from logit normalisation, again pointing towards LS. Given our analysis and empirical results, we strongly recommend logit normalisation for LS-trained models when performing SC. On the other hand, Fig. 5 also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE model. This explains why for CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.
@@ 6 Related work
Prediction with rejection. Selective classification falls into the broader problem setting of prediction with rejection. In the case of SC, misclassifications are to be rejected (El-Yaniv & Wiener, 2010). The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020a; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023). Another scenario is out-of-distribution (OOD) detection (Yang et al., 2021; Hendrycks & Gimpel, 2017), where data from outside of the training distribution are to be rejected. There is a plethora of research in this field (Xia & Bouganis, 2022a; Sun et al., 2021; Liu et al., 2020; Zhang et al., 2023; Liu et al., 2023; Wang et al., 2022; Hendrycks et al., 2022). Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data (Jaeger et al., 2023; Xia & Bouganis, 2022b; Kim et al., 2023). Notably, Deep Ensembles (Lakshminarayanan et al., 2020) have arisen as a reliable method to improve performance in all three scenarios (Kim et al., 2023; Xia & Bouganis, 2023; Laurent et al., 2023; 2024). We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
@@ 7 Concluding remarks
In this work, we elucidate the effect of label smoothing (LS) on selective classification (SC). Our experiments across various tasks and architectures show that LS leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves accuracy. By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions. We then investigate post-hoc logit normalisation as a method to improve the degraded SC performance caused by LS. We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced suppression by increasing the uncertainty when the max logit is higher. We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation. A further discussion about the practical impact of our work, as well as potential future work can be found in Appendix H.
Acknowledgments
Guoxuan Xia’s PhD is funded jointly by Arm and the EPSRC. This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-[AD011011970R3]). We thank the anonymous reviewers for their valuable feedback.
@@ Acknowledgments
To ensure transparency, we use publicly available datasets: CIFAR-10/100, ImageNet, Cityscapes, IMDB, Bank Marketing, and Online shoppers. Our detailed experimental methods are outlined in Appendix C. Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix E.
To help replicate our work, we share our source code on GitHub based on TorchUncertainty (Lafage et al., 2025), [^7] https://github.com/ENSTA-U2IS-AI/Label-smoothing-Selective-classification-Code, including the configuration files and training code. We also release the most important models on Hugging Face. [^8] https://huggingface.co/ENSTA-U2IS/Label-smoothing-Selective-classification
@@
Appendix A Glossary of Notation
We summarise the main notations used in the paper in Tab. 1.
Table 1: Glossary of Notation
Notation Meaning $p(\cdot)$ Probability density function $P(\cdot)$ Probability mass function $\boldsymbol{x}$ Input datum in $\mathbb{R}^{D}$ $y$ Label (categorical) from the set of $K$ possible labels $\{\omega_{k}\}_{k=1}^{K}$ $p_{\text{data}}(\cdot)$ True data distribution $\mathcal{D}_{\text{tr}}=\{\boldsymbol{x}^{(n)},y^{(n)}\}_{n=1}^{N}$ Training dataset of inputs and labels drawn from $p_{\text{data}}(\boldsymbol{x},y)$ $\boldsymbol{\theta}$ Model parameters $\boldsymbol{v}$ Logits (pre-softmax model outputs) in $\mathbb{R}^{K}$ $v_{\text{max}}$ Maximum logit, $\max\nolimits_{k}v_{k}(\boldsymbol{x};\boldsymbol{\theta})$ $\bar{\boldsymbol{\pi}}(\boldsymbol{x})$ True categorical conditional data distribution, $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})$ Softmax output in $[0,1]^{K}$ of model, $\pi_{k}=\frac{\exp v_{k}}{\sum_{i}\exp v_{i}}=P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})$ ()(-)( - )MSP/$\pi_{\text{max}}$ Maximum softmax probability, $\max\nolimits_{k}\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})$, negate for uncertainty $\delta_{ij}$ Kronecker delta, $\delta_{ij}=1\text{ if }i=j,\text{ and }0\text{ if }i\neq j$ KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] Kullback–Leibler divergence $\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ Empirical cross entropy loss minimised over finite data $\mathcal{D}_{\text{tr}}$ $\mathcal{L}_{\text{CE}}^{\text{true}}(\boldsymbol{\theta})$ True cross entropy loss minimised over $p_{\text{data}}(\boldsymbol{x},y)$. $\alpha$ Label smoothing parameter in $[0,1]$ $\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ Label smoothing loss $\boldsymbol{x}^{*}$ Test input datum $f(\boldsymbol{x})$ Classifier function $\hat{y}$ Label predicted by model for a given input, $\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}\nolimits_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})$ $\bar{\pi}_{\hat{y}}$ True probability of label predicted by model $\bar{\pi}_{\hat{y}}=P_{\text{data}}(\hat{y}|\boldsymbol{x})$ $P_{\text{error}}$ Probability of a given label prediction being incorrect, $P_{\text{error}}=1-\bar{\pi}_{\hat{y}}$ $U(\boldsymbol{x})$ Scalar uncertainty score $g(\boldsymbol{x};\tau)$ Binary rejection function, 0 (reject) if $U>\tau$ otherwise 1 (accept) $\text{Risk}(f,g;\tau)$ Average error on accepted samples for threshold $\tau$ $\text{Coverage}(g;\tau)$ Proportion of all data that is accepted for threshold $\tau$ $\frac{\partial\mathcal{L}_{\text{sup}}}{\partial v_{k}}$ Suppression logit gradient, difference between LS and CE gradients, $\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}$ $\boldsymbol{v}^{\prime}$ $p$-normalised logits, $\boldsymbol{v}^{\prime}=[\boldsymbol{v}+s]/\left\lVert\boldsymbol{v}+s\right\rVert_{p}$ with scalar shift $s$.
Appendix B Additional results
B.1 Illustrative examples of ImageNet overconfidence
@@
B.2 Complete results for ViT and DeepLabV3+
@@
  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix G.
Figure 10: RC curves evaluating on the combination of ImageNet and ImageNet-sketch. For low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced. This shows that LS leads to increasing overconfidence on less-well-fit data.
@@
CE #samples 7940 1148
#errors 94 74
error rate 1.2 6.4
LS $\alpha=0.1$samples 7422 1666
#errors 174 296
error rate 2.3 17.8
LS $\alpha=0.2$samples 6808 2280
#errors 225 549
error rate 3.3 24.1
LS $\alpha=0.3$samples 6726 2362
#errors 258 647
error rate 3.8 27.4
@@
CE #samples 8529 559
#errors 90 35
error rate 1.1 6.3
LS $\alpha=0.1$samples 8229 859
#errors 158 102
error rate 1.9 11.9
LS $\alpha=0.2$samples 7927 1161
#errors 196 161
error rate 2.5 13.9
LS $\alpha=0.3$samples 7618 1470
#errors 254 294
error rate 3.3 20.0
@@
B.3 Small-scale CIFAR experiments
@@
B.4 Small-scale tabular data experiments
To increase the scope of our experiments on the degradation of selective classification due to label-smoothing, we perform small-scale tabular binary classification. Following the setting of TabTransformer (Huang et al., 2020b), we train two-hidden-layer MLPs on Bank Marketing (Moro et al., 2014) and Online purchasing shoppers intentions (Sakar et al., 2019).
Figure 17: RC curves for MLPs trained on Bank Marketing and Online Shoppers using $\alpha=0.6$, hence with a gap between the positive and negative smooth labels of 0.4.
In this binary classification setting, we set $\alpha$ to 0.6 to sufficiently reduce the gap between the optimal MSP corresponding to the positive and negative detections of the class at stake. With $\alpha=0.6$ the minimum of the loss with label-smoothing is achieved for a prediction with a score of 0.7 when the hard label equals 1, and a score of 0.3 when the hard label is 0, resulting in a gap of a magnitude of 0.4. In Fig. 17, we see that in both cases, the AURC of the LS models are worse than those trained with the classical cross-entropy despite their accuracy being very similar. Furthermore, we see that the CE models are able to get much closer to zero risk, similar to the image classification experiments.
However, we note that the effect of label smoothing seems less pronounced for small-scale tabular data. We posit that this may be due to the data distributions being simpler and easier to capture, such that the neural networks are generally better fit and more knowledgeable about the data. Thus reducing the level of imbalanced suppression over training data (see Sec. 4). To address this lack of difficulty in fitting the distribution, we reduce the number of training points (as described in Appendix C).
Moreover, training small-scale and simple datasets drastically reduces the quality of our estimation of the Risk-Coverage curves. Given the limited number of data and the very high accuracy achieved by our MLPs, the number of errors used to estimate the risk is limited. The estimation of the “true” risk-coverage curve – which would be obtained with the whole distribution – is therefore imprecise and suffers from an important variability. The results may thus differ when starting the optimization process from different initializations, using a different batch composition and order, or due to the non-deterministic nature of some algorithms used to compute the backpropagation (Laurent et al., 2023).
B.5 Small-scale natural language data
With also provide small-scale natural language processing experiments training LSTM-based models (Sepp Hochreiter, Jürgen Schmidhuber, 1997) combined with two-layer perceptrons. We focus on LSTMs to add another architecture, and given that CNNs and transformers were already used in the image-classification setting (see e.g. Sec. 4).
Figure 18: RC curves for LSTMs trained the IMDB Movie review dataset, using $\alpha=0.6$, hence with a gap between the positive and negative smooth labels of 0.4.
We train two networks on the IMDB Movie Review dataset (Maas et al., 2011) with the classical binary cross-entropy loss and our implementation of the binary cross entropy with label-smoothing. Similarly to our results in image classification and tabular data classification, Fig. 18 shows that the LSTMs trained with label smoothing have worse AURC than models trained with cross-entropy despite their lower error-rate. Furthermore, the models also display greater error rates at high-confidence (low-coverage). We provide more extensive details on the training in Appendix C.
Appendix C Reproducibility
[…]
Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix B. All our models are trained with PyTorch (Paszke et al., 2019). We use the native implementation from the CrossEntropyLoss for label smoothing in the multi-class setting and a custom version of the BCEWithLogitsLoss for the binary experiments as the original does not support label smoothing. We use the original implementation of the authors for the negative label smoothing (Wei et al., 2022) experiments.
NB: The results presented in the final camera-ready ICLR 2025 version of this paper may differ slightly to earlier preprint versions. This is due to models being retrained with different random seeds, train-val-test splits, and having minor recipe and hardware differences after a consolidation of the experimental codebase to TorchUncertainty. [^11] https://torch-uncertainty.github.io/ Readers should use the information in this final version for reproducibility purposes.
C.1 Image classification